# Speak the Same Language: Global LiDAR Registration on BIM Using Pose Hough Transform

### v1

@@ VI-B Performance on Registration Recall
The registration recall results on the LiBIM-UST dataset are summarized in Table III. This table compares the performance of image registration, point cloud registration, and our method.
Methods relying on local descriptors, such as SIFT [31], Superpoint [32], and FPFH [34], all recorded a 0% registration recall. Despite attempts to optimize these methods by adjusting descriptor parameters and increasing the number of correspondences (typical counts were 400 for SIFT and Superpoint, and 4000 for FPFH), none were successful. The failure of these methods can be primarily attributed to two factors:
  • Inherent Limitations of Local Descriptors: The LiDAR submap and BIM share only basic geometric structures such as walls and corners, which often exhibit repetitive patterns (like single planes or simple intersections). This leads to local descriptors that are not discriminative enough for effective registration.
  • Mismatch with BIM’s Geometric Information: For image descriptors, the lack of rich gradient information due to uniform heights of indoor floors and ceilings poses a challenge. Point cloud descriptors like FPFH struggle because they require a diverse distribution of normal vectors, which is not typical in the mostly parallel or perpendicular planes found in BIM environments.
The 3D-BBS method [36] showed potential for LiDAR to BIM registration by searching through transformation space to align submaps closely with BIM. However, it does not account for discrepancies between as-designed and as-built conditions, and the extensive size of BIM compared to submaps makes the search process for 3D-BBS quite time-consuming.
Our method outperforms the aforementioned techniques by leveraging only the geometric features common to both the LiDAR submap and BIM, such as corners and walls. This approach not only reduces computational overhead but also minimizes interference from non-shared structures. By employing a Hough transform coupled with a voting mechanism, our method ensures that the pool of transformation candidates maximally includes potential true transformations, thereby enhancing the registration recall. Additionally, unlike the exhaustive search in 3D-BBS, our use of triangle descriptor-based matching significantly narrows the search scope. We also observed that a larger submap size ($d_{s}$) tends to increase recall, indicating its benefit in preventing registration degeneration, particularly in corridor settings.
@@
Variable Value Registration Recall(%) Average Time(s)
Building Day 2f-office Building Day 2f-office
Triangle Descriptor $r_{\mathrm{l}}$ 0.5m 83.33 73.33 15.96 30.17
1.25m 91.67 93.33 25.44 158.16
2.0m 86.11 73.33 39.16 236.20
Triangle Descriptor $r_{\mathrm{a}}$ 2 88.89 80.00 10.34 57.47
3.8 91.67 93.33 25.44 158.16
4.5 88.89 80.00 25.58 145.08
Hough Transform $r_{\mathrm{xy}}$ 1.5m 91.67 86.67 28.95 161.41
2.5m 91.67 93.33 25.44 158.16
4.0 m 91.67 60.00 18.05 121.85
Hough Transform $r_{\mathrm{yaw}}$ 3 91.67 93.33 28.59 159.80
5 91.67 93.33 25.44 158.16
7 91.67 86.67 22.61 124.59
Top J in Voting 1 86.11 46.67 14.40 156.55
5 91.67 66.67 19.01 156.84
10 91.67 80.00 23.75 159.82
15 91.67 93.33 25.44 158.16
20 91.67 93.33 27.92 160.54
Transformation Estimator RANSAC 38.89 20.00 24.69 153.42
TEASER++ 77.78 26.67 NA NA
Hough Voting 91.67 93.33 25.44 158.16
Figure 8: The registration recall and average time with varing parameter $J$ in voting process.
@@ VI-C Ablation Study
In this section, we explore the effects of several key parameters on the overall performance of our registration system. These parameters include the side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ for triangle descriptors, the x-y resolution $r_{\mathrm{xy}}$, yaw resolution $r_{\mathrm{yaw}}$, and the parameter $J$ in the voting and verification process, as well as different back-end systems for pose estimation. We set $d_{s}=10m$ for Building Day sequences and $d_{s}=30m$ for office sequences by default.
@@ VI-C1 Triangle descriptor
The side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ are crucial in addressing the discrepancies between corners derived from submaps and those from BIM. These discrepancies typically arise from LiDAR odometry drifts, differences between the as-designed BIM and the as-built environment, and errors in the corner extraction algorithm. As shown in Figure 9 and Table IV, both excessively high and low resolutions detrimentally affect the registration recall. Excessively fine resolutions prevent correct corner triplet correspondences from being grouped under the same triangle descriptor, leading to increased false negatives. On the other hand, excessively coarse resolutions may group incorrect correspondences together, increasing false positives and introducing incorrect transformation candidates during the voting process. This not only complicates the transformation verification process but also increases computation time significantly due to the higher number of corner triplets sharing the same descriptor, especially with an all-to-all matching strategy (see Section IV-C).
VI-C2 Transformation voting and verification
Table VI-C and Figure 8 illustrate how $r_{\mathrm{xy}}$, $r_{\mathrm{yaw}}$, and $J$ influence registration recall and computation time. Incorrect corner extraction introduces noise in the parameter space, complicating the voting process. Too fine a resolution scatters votes, preventing them from aggregating around the true transformation, while too coarse a resolution reduces the precision of estimated transformations and may lead to false positives in transformation verification. Adjusting $J$ affects both registration success and processing time; a smaller $J$ might overlook true transformations, particularly in structurally repetitive environments like office corridors, whereas a larger $J$ enhances recall at the cost of increased computation.
VI-C3 Back-end
As outlined in Section VI-A3, our framework incorporates the back-end algorithms, RANSAC and TEASER++, commonly employed in global localization methods such as STD [28] and Outram [23]. Specifically, for each subBIM candidate, we apply RANSAC and TEASER++ to generate a transformation candidate from its corner-wise correspondences. Then, we utilize the proposed wall-based verification to find the optimal candidate.
Table IV presents the registration results following the integration of RANSAC and TEASER++ into our system. Given that each subBIM candidate produces only a single transformation candidate, the likelihood of overlooking the true transformation is heightened. Furthermore, with a high outier ratio, RANSAC needs exponential time complexity to find the true estimation. TEASER++ trades off spatial complexity for speed with an $O(n^{2})$ spatial complexity, and constructing the compatibility graph also demands a time complexity of $O(n^{2})$, where $n$ is the number of corner-wise correspondences. This complexity has occasionally led to memory overflow during execution, causing termination of the program on some submaps. Consequently, in such instances, the runtime for TEASER++ is not recorded and is denoted as NA.
Figure 9: The registration recall and average time with varing side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ in the construction of triangle descriptors.
Figure 10: Global registration results using our our method. Unsuccessful cases are illustrated in Failure-1 and Failure-2. Failure-1: multiple highly similar patterns exist. Failure-2: the scenario is geometrically uninformative like a long and narrow corridor, where only parallel lines can be extracted in the submap as shown in the dashed box.
VI-D Case Studies
This subsection highlights specific scenarios that demonstrate the limitations of our proposed method. As depicted in Figure 10, our method struggles in environments characterized by repetitive architectural patterns which lack distinctive corner triplets. Additionally, in scenarios where the mobile platform navigates through narrow and elongated corridors, the observable effective corner points are markedly reduced. This scarcity of corner points often leads to what we term as ‘degeneration issues,’ where the lack of sufficient distinctive features compromises the algorithm’s performance.
Figure 10 also provides further visualizations of the registration outcomes across different scenarios, supplemented with visual images captured by fisheye cameras mounted on the mobile platform. These visual aids illustrate the practical challenges faced in these specific architectural settings and help in understanding the limitations of the current system in managing complex environments.
VII Conclusion
In this study, we design a global registration method that could align LiDAR data to BIM without initial guess. Lines and corners are extracted in BIM and LiDAR as shared features. Then, triangle descriptors are proposed and correspondences are formed using the Hash table. In the back end, we incorporate the Hough transform to estimate the multiple candidates for transformation. Lastly, we verify the consistency of each candidate and find the optimal transformation as the final pose estimation. The method is evaluated on a real-world university building, with collected multi-session LiDAR data, named LiBIM-UST dataset.
@@ VII Conclusion
We sincerely thank Prof. Jack Chin Pang Cheng, a Professor at the Department of Civil and Environmental Engineering, Hong Kong University of Science and Technology, for kindly providing the BIM of the HKUST Academic Building.

### v2

@@ VI-B Performance on Registration Recall
The registration recall results on the LiBIM-UST dataset are summarized in Table III. This table compares the performance of image-based registration, point cloud-based registration, and our method.
Methods relying on local descriptors, such as SIFT [34], Superpoint [35], and FPFH [37], all recorded a 0% registration recall. Despite attempts to optimize these methods by adjusting descriptor parameters and increasing the number of correspondences (typical counts were 400 for SIFT and Superpoint, and 4000 for FPFH), none were successful. Similarly, the learning-based Geotransformer [22] suffers from the cross-modality challenges and result in an extremely low registration recall. The collective failure of local descriptor-based methods highlights the ineffectiveness of this approach for LiDAR and BIM registration. This is because these methods typically rely on discriminative local features, such as normals and gradients, which cannot be adequately provided by BIM. BIM structures predominantly consist of planar surfaces, where normals are mostly parallel or perpendicular, offering limited discriminative power. In contrast, our proposed triangle descriptor avoids relying on these constrained local features and instead harnesses more distinctive and shared structural information, enabling more robust data association.
The 3D-BBS method [39] showed potential for LiDAR-to-BIM registration by searching through the transformation space to align submaps closely with the BIM. However, it does not account for discrepancies between as-designed and as-built conditions, leading to lower recall on most sequences. Additionally, since it searches the entire transformation space, 3D-BBS is computationally expensive and time-consuming. In contrast, our triangle descriptor-based matching drastically narrows the search scope. Furthermore, we introduce a hierarchical voting mechanism to accelerate the search process, significantly reducing computation time. The occupancy-aware score design further improves robustness to construction discrepancies, enabling a more reliable evaluation of alignment quality for each transformation candidate.
Figure 8: (a) and (b): The registration recall and average time with varying side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ for building triangle descriptors. (c) and (d): The registration recall and average time with varying translation resolution $r_{\mathrm{xy}}$ and yaw resolution $r_{\mathrm{yaw}}$ in the Hough transform.
@@
Variable Value Registration Recall(%) Average Time(ms)
Triangle Descriptor $r_{\mathrm{l}}$ 0.3m 87.60 88.29
0.5m 89.15 109.01
0.7m 88.28 152.38
Triangle Descriptor $r_{\mathrm{a}}$ 1 86.44 77.80
3 89.15 109.01
5 88.38 183.02
Hough Transform $r_{\mathrm{xy}}$ 0.10m 88.57 108.44
0.15m 89.15 109.01
0.20m 88.38 109.37
Hough Transform $r_{\mathrm{yaw}}$ 0.5 87.89 110.22
1.0 89.15 109.01
1.5 88.48 110.06
Hierachical Voting Vanilla 38.92 75.99
Layer 1 89.64 134.26
Layer 2 89.16 166.94
all 89.15 109.01
Transformation Estimator SVD 0 55.45
RANSAC 8.03 24027.81
TEASER++ 25.07 13851.75
Hough Voting 89.15 109.01
@@ VI-C Ablation Study
In this section, we explore the effects of several key parameters on the overall performance of our registration system. The parameters under evaluation include the building of the triangle descriptor, the Hough transform, hierarchical voting, and the transformation estimator. All sequences are considered for evaluation.
@@ VI-C1 Triangle descriptor
The side length resolution $r_{\mathrm{s}}$ and angle resolution $r_{\mathrm{a}}$ account for inaccuracies in corner extraction from the submaps and the differences between the as-designed BIM and the as-built environment, as the triangle descriptors do not need to be identical; they can still be matched as long as they fall into the same hash key.
@@
As shown in Figure 8-8 and Table IV, both excessively high and low resolutions adversely affect registration recall. Resolutions that are excessively fine prevent correct corner triplet correspondences from being grouped together, leading to an increase in false negatives. Conversely, excessively coarse resolutions may group incorrect correspondences together, increasing false positives and introducing incorrect transformation candidates during the voting process. This not only complicates the transformation verification process but also significantly increases computation time due to the higher number of corner triplets sharing the same descriptor, especially with an all-to-all matching strategy (see Section IV-C).
VI-C2 Hough Transform
The use of $r_{\mathrm{xy}}$ is analogous to $r_{\mathrm{s}}$ and $r_{\mathrm{a}}$, as both discretize the space into voxels to group similar values together. Resolutions that are too fine scatter the votes, preventing them from converging around the true transformation. Conversely, excessively coarse resolutions reduce the precision of the estimated transformations and may introduce false positives during transformation verification.
However, the Hough transform is less sensitive to resolution changes than the triangle descriptor construction, as the hierarchical voting mechanism effectively filters out false candidates, as shown in Figure 8 and 8. Additionally, the yaw resolution is less critical than the translation resolution, as our method primarily focuses on wall alignment, where yaw candidates are typically multiples of 90 degrees.
VI-C3 Hierarchical Voting
To compare the proposed hierarchical voting method with the original voting method, we evaluate our hierarchical scheme by retaining only the first $n$ layers, such as Layer 1 and Layer 2, where $n=1$ and $n=2$, with the default parameters $L=10000$ and $K=5000$, respectively. Additionally, we implement the original voting method, referred to as vanilla voting, which selects the transformation with the highest number of votes as the final result.
As shown in Table IV, Layer 1 achieves the best recall performance because it retains the largest number of transformation candidates. However, its runtime is the longest due to the computational burden of the verification step. On the other hand, vanilla voting exhibits the worst performance, as relying solely on the vote count as a criterion is limited by the insufficient information available from corner matching, as discussed in Section V-B. Conversely, the proposed hierarchical voting method achieves an optimal balance between efficiency and robustness.
VI-C4 Back-end Estimation
We compare the proposed transformation estimator with other alternative estimators, such as ordinary least squares using SVD [31], the robust estimator TEASER++ [38], and RANSAC [40]. As shown in Table IV, the proposed method achieves the best performance in terms of registration recall and average time consumption. SVD is highly sensitive to outliers and noise, resulting in a lower recall rate. Although TEASER++ is robust to outliers, it provides only a single transformation candidate, which may overlook the true transformation. RANSAC employs a hypothesis-and-verification strategy; however, the score of each candidate relies solely on consistent corner correspondences, which may be insufficient to comprehensively evaluate the transformation quality.
In terms of computational complexity, the proposed method achieves $O(N)$ complexity for the Hough transform and $O(N\log(N))$ for hierarchical voting. In comparison, alternative methods such as TEASER++ [38] have a worst-case complexity of $O(n^{2})$, where $N$ and $n$ denote the number of corner triplet correspondences and the number of corner correspondences, respectively, with $n<N\ll n^{2}$. Similarly, RANSAC-based methods [40] can exhibit exponential complexity under high outlier ratios. Furthermore, since calculating transformations from corner triplet correspondences and evaluating their confidence scores are independent processes, they can be executed in parallel. This parallelization significantly accelerates real-world implementation.
Figure 9: Precision and recall curves of the proposed occupancy-aware confidence score compared with other variants and methods. OSC denotes the proposed confidence score. OSC-1 denotes the score considering only the award component. OSC-2 includes an additional award component where both the submap and BIM are free. OSC-3 introduces an additional penalty component where the submap is occupied but the BIM is free.
VI-D Registration Reliability Assessment
The proposed occupancy-aware confidence score is designed to assess the reliability of registration. To evaluate its effectiveness, we construct a new benchmark in addition to the LiBIM-UST dataset. This benchmark includes unregistrable pairs created by combining cross-floor submaps and BIM models. The confidence score is employed to determine whether the estimated transformation is correct. For instance, if the score is high for a correctly registered pair of submaps and BIMs, it is considered a true positive. Conversely, if the score is high for an unregistrable pair, it is classified as a false positive. To analyze its performance, we plot the precision-recall (PR) curves in Figure 9.
The compared methods include the proposed occupancy-aware confidence score, the Chamfer distance (with distances truncated at 1 m), the overlap ratio [41] (with a neighborhood radius of 1 m), and the entropy-based distance [42] (with a neighborhood radius of 1 m). Additionally, we consider two specific cases within the confidence score: (1) scenarios where both the submap and BIM are free, treated as an award component, and (2) scenarios where the submap is occupied but the BIM is free, treated as a penalty component. These cases, which were not included in the original score, are detailed in Section V-B. Moreover, we also evaluate the impact of removing the original penalty component. The results demonstrate that the proposed confidence score (OSC in Figure 9) effectively distinguishes between correct and incorrect registrations, outperforming the compared methods.
Figure 10: Featured scenarios of LiDAR submap-to-BIM registration are presented, including four success cases and one failure case. In each case, the following are shown: RGB images; the submap, highlighting matched walls in green and unmatched walls in red; the BIM with matched walls, corners, and triplets, and the registration result along with its confidence score. Two special types, construction inconsistencies and unstructured walls, are highlighted in yellow and brown, respectively. In Success-1, a jagged wall, marked by a brown circle, is identified as a plane and successfully matched. A yellow cross indicates an extra bookshelf, identified as an unmatched plane. In Success-2, the top yellow cross indicates a misaligned wall caused by a large slanted LED display that exists in the actual environment but is not present in the BIM model. The bottom yellow crosses indicate lockers on both sides of a corridor, which visually narrow the original corridor. In Success-3, brown circles highlight two curved walls that are segmented into several small planes and successfully matched. In Success-4, the yellow cross indicates a transparent glass partitioned in the real world, which corresponds to an unmatched solid wall in the BIM. The failure case, Failure-1, shows an ambiguous long corridor scenario, where even the ground truth transformation yields a lower confidence score.
VI-E Case Studies and Limitation Discussion
VI-E1 Overview of Case Studies
Figure 10 illustrates several cases of LiDAR submap-to-BIM registration. To clarify the registration process, RGB images, extracted wall and corner features, constructed corner triplets, and the proposed confidence score are presented. Two special cases, construction inconsistencies and unstructured walls, are highlighted in yellow and brown, respectively.
VI-E2 Construction Inconsistencies
As shown in Figure 10, construction inconsistencies typically manifest in three forms: unconstructed parts (i.e., elements present in the BIM model but missing in reality, such as the transparent glass in Success-4), additional constructed parts (i.e., elements absent in the BIM model but present in reality, such as the bookshelf in Success-1 and lockers in Success-2), and position biases (i.e., elements that exist in both but are misaligned, such as the top image in Success-2). While these inconsistencies increase the number of unmatched triplets, our method prioritizes finding a sufficient number of matched triplets rather than ensuring a perfect bijective correspondence between corners. Theoretically, we only need to construct at least one correct corner triplet correspondence for pose estimation, while the proposed confidence score ensures the identification of the correct transformation candidate during the verification stage.
Furthermore, in computing the confidence score, scenarios where the submap is occupied but the BIM is free are not considered. This design ensures that while such inconsistencies, like additional constructed parts, do not contribute positively to the registration, they also do not negatively affect it in most cases. In summary, construction inconsistencies may impact the number of matched corner triplets and the quality of the confidence score. However, as long as the true transformation candidate outperforms others during the verification step, successful registration can still be achieved.
VI-E3 Unstructured Walls
Unstructured walls are represented as jagged walls (e.g., in Success-1) and curved walls (e.g., in Success-3) in the provided examples. Thanks to the plane threshold parameter $\sigma_{\lambda}$, our method does not require walls to be perfectly planar. Jagged walls are successfully detected as planes and matched accordingly. Curved walls are segmented into several small wall sections, which may produce several inaccurate corners. However, due to the design of our approach, which accounts for handling inaccurate corners through the discretization of descriptor and transformation spaces, curved walls still contribute positively to the registration process.
VI-E4 Limitations of the Proposed Method
While the above cases demonstrate the robustness of our method in handling challenging scenarios, there are still several limitations. First, our method requires at least one matched corner triplet and a sufficient number of matched walls to achieve a high confidence score. This means that larger submaps with more walls are more likely to result in successful registration. Second, although our method handles jagged and moderately curved walls effectively, it struggles with extreme cases, such as large curved walls, which fail to produce stable corners. Third, as illustrated in Failure-1, our method is affected by long corridor scenarios due to perceptual aliasing, a well-known systematic issue where two different locations exhibit similar appearances. To address this, additional information is needed to resolve ambiguities, such as enlarging the submap size. Alternatively, incorporating visual information could provide semantic cues, such as construction materials, functional zoning, or room divisions, which can be matched with BIM semantics to eliminate ambiguities.
VII Conclusion and Future Work
In this study, we propose a one-shot global registration method capable of aligning LiDAR data to BIM without requiring an initial guess. Walls and corners are extracted from both BIM and LiDAR data as shared features. Triangle descriptors are then computed based on corner triplets and matched using a Hash table, enabling fast and efficient retrieval independent of the BIM model’s size. In the back end, we incorporate the Hough transform and a hierarchical voting mechanism to hypothesize multiple transformation candidates. Subsequently, we introduce an occupancy-aware score to identify the optimal pose estimation. In the LiBIM-UST dataset, we demonstrate the effectiveness of our method in comparison with traditional image and point cloud registration techniques.
Future research will encompass both robotics and construction automation. In robotics, integrating global localization methods, such as the particle filter, presents a promising direction for enhancing applications that utilize BIM as a navigational map. We also plan to leverage the semantic information of BIM to enhance the performance. For construction automation, achieving global alignment between point clouds and BIM is not enough; it is also critical to identify dissimilar locations and elements to enable BIM correction for quality management. This capability is crucial for supporting AR applications on BIM.
@@ VII Conclusion and Future Work
We sincerely thank Prof. Jack Chin Pang Cheng for providing the BIM of the HKUST Academic Building.